OpenStudy (anonymous):

Write the equation of the circle whose center is at the origin and whose diameter is 12.

OpenStudy (anonymous):

x2+y2=sqrt 6

OpenStudy (hoblos):

well it is x2 + y2 = 6^2

OpenStudy (anonymous):

yah mistyped that one !!